### v2

# MAGPIE: Multi-Task Analysis of Media-Bias Generalization with Pre-Trained Identification of Expressions

Abstract
Media bias detection poses a complex, multifaceted problem traditionally tackled using single-task models and small in-domain datasets, consequently lacking generalizability. To address this, we introduce MAGPIE, the first large-scale multi-task pre-training approach explicitly tailored for media bias detection. To enable pre-training at scale, we present Large Bias Mixture (LBM), a compilation of 59 bias-related tasks. MAGPIE outperforms previous approaches in media bias detection on the Bias Annotation By Experts (BABE) dataset, with a relative improvement of 3.3% F1-score. MAGPIE also performs better than previous models on 5 out of 8 tasks in the Media Bias Identification Benchmark (MBIB). Using a RoBERTa encoder, MAGPIE needs only 15% of finetuning steps compared to single-task approaches. Our evaluation shows, for instance, that tasks like sentiment and emotionality boost all learning, all tasks enhance fake news detection, and scaling tasks leads to the best results. MAGPIE confirms that MTL is a promising approach for addressing media bias detection, enhancing the accuracy and efficiency of existing models. Furthermore, LBM is the first available resource collection focused on media bias MTL.
Keywords: Media bias, Multi-task learning, Text classification
[…]
MAGPIE: Multi-Task Analysis of Media-Bias Generalization
with Pre-Trained Identification of Expressions
[…]
${}^{1}$zech Technical University, Prague, Czech Republic
${}^{2}$niversity of Göttingen, Göttingen, Germany, {last}@uni-goettingen.de
${}^{3}$ational Institute of Informatics, Tokyo, Japan, {last}@nii.ac.jp
${}^{4}$, TU Munich, Germany, ${}^{5}$niversity of Konstanz, Germany, {last}@uni-konstanz.de
${}^{6}${t.horych, m.wessel, t.spinde}@media-bias-research.org
Abstract content
1.   Introduction
Media bias is the skewed portrayal of information favoring certain group interests Recasens et al. (2013), which manifests in multiple facets, including political, gender, racial, and linguistic biases. Such subtypes of bias, which can intersect and coexist in complex combinations, make the classification of media bias a challenging task Raza et al. (2022). Existing research on media bias detection primarily involves training classifiers on small in-domain datasets Krieger et al. (2022), which exhibit limited generalizability across diverse domains Wessel et al. (2023).
This paper builds upon the work of Wessel et al. (2023), emphasizing that the multifaceted nature of media bias detection requires a shift from isolated approaches to multi-task methodologies, considering a broad spectrum of bias types and datasets. The recent advancements in Multi-Task Learning (MTL) Aribandi et al. (2021a); Chen et al. (2021); Kirstein et al. (2022) open up promising opportunities to overcome these challenges by enabling knowledge transfer across domains and tasks. Despite the potential, a comprehensive MTL approach for media bias detection is yet to be realized. The only other media bias MTL method Spinde et al. (2022) underperforms due to its narrow task focus and does not surpass baseline outcomes.
Figure 1: MAGPIE has a pre-trained representation of multiple biases (persuasive, subjective, etc.). This enables it to outperform models based on single-task learning (STL) paradigms.
In this study, we make five main contributions:
  • We present MAGPIE (Multi-Task Media-Bias Analysis Generalization for Pre-Trained Identification of Expressions) the first large-scale multi-task pre-training approach for media bias detection. By including diverse biases like persuasive and subjective, a classifier trained on top of MAGPIE correctly classifies sentences that state-of-the-art single-task models misidentify (we show an example in Figure 1).
  • We introduce LBM (Large Bias Mixture), a pre-training composition of 59 bias-related tasks encompassing wide range of biases such as linguistic bias, gender bias and group bias.
  • We provide an analysis of a task selection and demonstrate the effectiveness of scaling the number of tasks.
  • We demonstrate that MAGPIE outperforms the previous state-of-the-art model by 3.3% on the Media Bias Annotation by Experts (BABE) dataset Spinde et al. (2021c) and achieves state-of-the-art results on the Media Bias Identification Benchmark (MBIB) collection Wessel et al. (2023).
  • We make all resources, including datasets, training framework, documentation, and models, publicly available on GitHub.
    github.com/magpie-multi-task
These contributions highlight the potential of MTL in improving media bias detection. Our findings show, e.g., that tasks like sentiment and emotionality enhance overall learning, all tasks boost fake news detection, and scaling tasks leads to optimal results. Another key insight of our research is the value of MTL in contexts where the primary dataset is small [^1] For example, the Media Bias Annotation by Experts (BABE) dataset Spinde et al. (2021c).. By learning generalized bias knowledge from a range of tasks, we can improve the accuracy and efficiency of existing models, even in the face of limited data. Overall, our research offers a novel multi-task learning approach with a variety of media bias tasks, contributing to the understanding of media bias detection.
2.   Related Work
2.1.   Media Bias
Media bias is a complex issue Lee et al. (2021a); Recasens et al. (2013); Raza et al. (2022) composed of varying definitions of bias subtypes such as linguistic bias, context bias, or group bias Wessel et al. (2023). In their literature review, Spinde et al. (2023) provide an extensive overview of research on media bias and related subtypes of bias.
Media bias detection approaches have evolved from hand-crafted features Recasens et al. (2013); Hube and Fetahu (2018); Spinde et al. (2021d) to neural models Spinde et al. (2022); Chen et al. (2021); Spinde et al. (2021c); Huguet Cabot et al. (2021); Sinha and Dasgupta (2021); Raza et al. (2022). However, existing models, so far, focus only on single tasks and saturate quickly on smaller datasets Wessel et al. (2023). As most neural approaches require large quantities of data, those relying on single and small datasets cannot provide a realistic scenario for their solutions (e.g., Fan et al. (2019)). We will first provide an overview of existing datasets and then show how to exploit their diversity within the media bias domain.
Media bias tasks and datasets mainly cover individual, self-contained tasks such as binary classifications Recasens et al. (2013); Spinde et al. (2021b), which, so far, are not explored in relation to each other Spinde et al. (2023). Wessel et al. (2023) systematically form the media bias detection benchmark MBIB by reviewing over 100 media bias datasets and consolidating 22 of them into eight distinct tasks like linguistic, racial, and political bias. Their study highlights that methods only focused on one of these tasks exhibit limitations in their detection capabilities. MAGPIE encompasses all the tasks identified in the MBIB but also significantly expands its scope by incorporating an additional 51 media bias-related tasks due to a variety of limitations in MBIB (see Section 3.1).
2.2.   Multi-Task Learning
MTL shows significant improvements in various NLP tasks, including sentiment analysis He et al. (2019), text summarization Kirstein et al. (2022), and natural language understanding Raffel et al. (2020). In MTL, a model leverages knowledge gained from one task to improve the performance of others. Aribandi et al. (2021b) demonstrate that increasing the number of tasks generally leads to improved performance for downstream NLP applications. Aghajanyan et al. (2021) show that pre-finetuning, a large-scale multi-task learning phase, consistently improves the performance and efficiency of pre-trained models across diverse tasks, with results improving linearly with the number of tasks beyond a certain threshold.
As described above, media bias can be seen as a composite problem composed of various interrelated bias types Wessel et al. (2023). In the realm of Natural Language Understanding (NLU), MTL has proven to be highly effective when incorporating related tasks Aribandi et al. (2021b). For instance, benchmarks such as GLUE and SuperGLUE successfully decompose the NLU problem into a suite of proxy tasks, including paraphrase detection and semantic evaluation, thereby substantially improving performance across a range of NLU tasks Wang et al. (2018, 2019). Motivated by this success in NLU, we propose to jointly learn from different bias types within the media bias domain. With this approach, we aim to treat media bias not as a singular entity but as many interconnected issues.
The selection of tasks is pivotal to the efficacy of MTL. There have been several attempts to automate task selection, including learning the data selection with Bayesian optimization Bingel and Søgaard (2017) or estimating task relations Ruder and Plank (2017). The most model-agnostic approach is GradTS Ma et al. (2021), which is highly scalable due to low resource requirements, and is therefore implemented within MAGPIE. GradTS accumulates gradients of attention heads and selects tasks based on their correlation with the primary task’s attention. The selected tasks are trained jointly and share representations across tasks.
3.   Methodolodgy
We implement MAGPIE using pre-finetuning as introduced in Aghajanyan et al. (2021) (See also Section 2). As such, MAGPIE is an encoder-only MTL transformer model pre-finetuned on 59 media bias-related auxiliary tasks provided by Large Bias Mixture (LBM), an extensive task collection of bias-related datasets [^2] LBM is also published within this paper, see Section 3.1.. We incorporate a novel approach of a Head-Specific Early Stopping and Resurrection to effectively handle tasks of varying sizes (Section 3.3.1).
Figure 2: The process of training and evaluating MAGPIE. The purple steps describe the construction and usage of LBM, the yellow the model training, and the green the model evaluation.
As outlined in Figure 2, our first step involves constructing the LBM. Following this, we define the model and multi-task learning (MTL) framework employed to train MAGPIE, which includes optimization strategies, task sampling, and auxiliary task selection. Lastly, we evaluate MAGPIE on two primary resources: the Media Bias Annotation by Experts (BABE) dataset [^3] BABE provides high-quality labels that capture a broad range of linguistic biases, thus allowing us to evaluate our model’s generalizability within a single dataset context. Spinde et al. (2021c), and the Media Bias Identification Benchmark (MBIB) collection [^4] MBIB provides a variety of bias tasks for testing and is currently the most extensive media bias benchmark available. Wessel et al. (2023).
3.1.   The LBM (Large Bias Mixture) task collection
Currently, MBIB is the only collection of media bias tasks. However, it does not include tasks that constitute a form of media bias [^5] As per the definition of directly and indirectly related media bias tasks given by Spinde et al. (2023). more indirectly, such as sentiment analysis or emotion classification. MAGPIE aims to integrate tasks both directly linked to media bias and those peripherally related, such as sentiment analysis, to provide broader coverage of linguistic features in the media bias context. Therefore, we introduce Large Bias Mixture (LBM), a more general collection of relevant media bias and media bias-related tasks, more suitable for our MTL approach. We show our task selection process in Figure 3.
First, we manually assess a list of 115 media bias-related datasets categorized into task families composed by Wessel et al. (2023). A task family is a conceptual grouping of tasks that share similar objectives, such as those related to gender bias, encompassing pronoun coreference resolution, gender classification, and sexism detection.
We use this notion of task families to analyze general knowledge transfer between media bias tasks, such as proposed by Aribandi et al. (2021b) for general language tasks.
We filter the collection of the 115 datasets based on the following criteria [^6] We base these on the criteria set by Wessel et al. (2023). However, we acknowledge that determining the dataset quality remains a manual and subjective choice.:
  • Accessibility: Datasets have to be publicly accessible.
  • Text granularity: We only use datasets labeled on a sentence level (not on, for instance, article or word level)
  • Quality of annotations: We exclude datasets with little to no documentation or unreliable annotation sources.
  • Duplicates: We filter out datasets that contain full or partial duplicates of each other
Of the 115 datasets collected, we discard 11 datasets that are not publicly available. We discard 52 with article-level annotations and 5 with annotations on other levels [^7] One dataset has annotations on only words, two on users, and three on outlets.. We remove 5 datasets due to low quality or unreliable annotations and discard 4 duplicates. Applying these criteria leaves 38 datasets. Including 8 handpicked datasets not originally listed gives us 46 datasets. These are categorized into task families ensuring no overlap and more than two datasets per family. Finally, ten datasets with multi-level annotations, e.g., token and sentence level, are split into tasks, yielding final number of 59 tasks.
Figure 3: The workflow of collecting datasets from the initial list to the final LBM collection.
In order to standardize examples from various domains, we apply a unified text-cleaning procedure to each dataset. The process involves: (1) discarding sentences with fewer than 20 characters, (2) removing URL links, special characters, and whitespace, and (3) eliminating duplicates potentially created with steps (1) and (2).
The final Large Bias Mixture (LBM) includes 59 tasks, categorized into 9 distinct task families, encompassing 1,210,084 labeled sentences. We make the LBM publicly accessible, to facilitate research in media bias detection and other computational-social-science tasks. References and short descriptions of all datasets and corresponding tasks can be found in Table 4.
3.2.   The Base Model
In terms of the base language model for our procedure, we adopt a pre-trained RoBERTa Liu et al. (2019) encoder due to its proven state-of-the-art performances across various media bias applications Spinde et al. (2021c); Krieger et al. (2022).
3.3.   The MTL framework
Pre-finetuning. To effectively harness the generalization abilities of MTL for media bias detection, we adopt a pre-finetuning procedure Aghajanyan et al. (2021). Pre-finetuning is a complementary approach to pre-training, where a model, already pre-trained on a typically unsupervised task, is subjected to another intermediate step of pre-training. While incorporating MTL directly into the pre-training stage has demonstrated performance gains Aribandi et al. (2021b), we opt for pre-finetuning as it offers significantly reduced computational demands while still capitalizing on the benefits of MTL Aghajanyan et al. (2021).
Sharing representations. We use hard parameter sharing to share the underlying encoder among all tasks while using separate task-specific heads. For each task, we attach dense layers, or "heads", to the shared encoder. These heads are optimized individually per task while the shared encoder learns general bias representations.
However, multi-task optimization presents challenges due to differing gradient directions and magnitudes Yu et al. (2020). For instance, two tasks, A and B, may have opposing gradients with the same magnitude, nullifying their sum. On the other hand, if Task A’s gradient greatly surpasses that of Task B, gradient A becomes dominant. We counter the gradient misalignment by using a variation of the PCGrad de-confliction algorithm and loss scaling Yu et al. (2020).
Conflicting gradients and loss scaling. In multi-task training involving $n$ tasks, encoder parameters receive $n$ potentially conflicting gradients. Efficient handling of this conflict, such as PCGrad Yu et al. (2020), requires storing a set of gradients for each task involved in the update, leading to infeasible memory requirements among our 59 LBM tasks. Therefore, we propose a variation of PCGrad we call PCGrad-online which preserves the fundamental idea of the original algorithm but is more memory efficient, requiring only one set of gradients instead of $n$ sets per update. Adopting Muppet’s method, we solve the issue of varying task gradient magnitudes by re-scaling the task loss with the inverse log size of its output space, ensuring balanced gradients and preventing task dominance in training steps Aghajanyan et al. (2021).
3.3.1.   Data sampling and early stopping
To prevent large tasks from dominating the optimization, we ensure uniform data distribution by sampling one fixed-size sub-batch from each task per training step, a regular approach in MTL Aribandi et al. (2021b); Spinde et al. (2022). We also employ early stopping as a regularization for each task individually to prevent over-fitting of tasks that converge faster. However, these methods often fall short when confronted with tasks of varied complexity and differing convergence speeds, which both is the case for tasks in LBM. When task A stops early while task B takes longer to converge, the latent representation of the shared encoder shifts toward task B. We aim to mitigate this issue by employing a training strategy that tackles the latent representation shift using two complementary approaches:
  1. 1.
    Head-Specific-Early-Stopping (HSES)
  2. 2.
    Resurrection
HSES. When the task stops, we stop updating its specific head parameters while still backpropagating its language model gradients. This method stems from the observation that not all tasks benefit from the shared layers’ continuous learning, especially after they have reached an optimal state.
Resurrection. When the task stops, we allow it to resume the training after its validation loss starts to increase again. This enables the task to adapt its head parameters to the latest latent representation.
HSES maintains quality of faster-converging tasks, while Resurrection allows further adaptation when needed. Their combination aims for balanced, adaptive learning for tasks with varied complexities and convergence speeds. We perform a preliminary evaluation of the effectiveness of this method in the in Section 4.4. However, we stress the need for further extensive analysis in Limitations.
3.3.2.   Auxiliary task selection
Multi-Task Learning often involves selecting well-used datasets, leading to potential selection biases. Furthermore, manually handpicking datasets becomes challenging due to varying and potentially ambiguous bias annotation schemes. To automate the process of task selection, we utilize the GradTS algorithm Ma et al. (2021). We choose GradTS due to its demonstrated efficiency and its simplicity of implementation, which enhances its usability.
Gradient-based Task selection. In line with GradTS, we construct auxiliary tasks as follows. We individually train all tasks, accumulate absolute gradients, then extract and layer-wise normalize these in the attention heads, forming a 12x12 importance matrix [^8] RoBERTa has 12 attention heads on each of the 12 layers. for each task. Tasks are sorted by correlation between each task’s matrix and BABE task’s matrix. We pre-finetune $m-1$ models on the first $k$ tasks from the sorted list, where $k$ varies from 1 to $m-1$ and $m$ is the size of LBM. The BABE task is then evaluated on these pre-finetuned models, with the optimal $k$ determined by evaluation loss. For further details on the GradTS algorithm, please see Ma et al. (2021).
3.3.3.   Hyperparameters and training stability
Finetuning transformer models on small datasets often leads to inconsistent outcomes, such as overfitting, training instabilities, or significant performance variabilities Dodge et al. (2020). To ensure reproducibility and consistency, we use a fixed random seed for each run and perform individual task hyperparameter grid searches. To negate the influence of random weight initializations, we evaluate and compare models using 30 random seeds, reporting average performance. We use $321$ as a random seed for all experiments. For the final evaluation, where we employ 30 random seeds, we use values from $0..29$. For optimizing the model, we use a per-task batch size of 32, an AdamW optimizer, and a polynomial learning scheduler. We run all experiments on 1 NVIDIA TESLA A100 GPU with 40 GB of memory.
4.   Empirical Results
In this section, we present the results of our MTL approach. First, we report the set of auxiliary tasks selected by GradTS for pre-finetuning. Next, we assess how the model pre-finetuned on the GradTS set performs during subsequent finetuning on the BABE task, compared to a random choice of tasks and a full set of LBM tasks. We also compare the MTL approach to a single-task baseline and multiple MTL baselines and evaluate the performance of our best model, MAGPIE, on the MBIB benchmark. Then, we analyze the LBM taxonomy through a study on knowledge transfer between families. Lastly, we evaluate the effects of the proposed methods, HSES and Resurrection, through a preliminary study.
4.1.   Auxiliary tasks selection
We select suitable auxiliary tasks by calculating the correlation between attention-head importance matrices. We use Kendall’s correlation coefficient, as suggested by Puth et al. (2015). We find a local minimum for the BABE evaluation loss when pre-finetuning on the first $k=10$ most correlated tasks. The final set of the ten most correlated tasks referred to as gradts set, is displayed in Table 1. The tasks in the gradts set demonstrate a strong semantic connection to media bias, encompassing areas such as lexical bias, rumor detection, and fake news detection.
Task Type $\tau$ (correlation)
Persuasive techniques 0.73
Lexical/Informational bias 0.72
Rumour detection 0.69
Sentiment analysis 0.68
Global warming stance detection 0.68
Subjective bias 0.67
Subjective bias 0.67
Veracity classification 0.64
Gender bias 0.64
Fake news detection 0.63
Table 1: Attention importance correlation wrt media bias task.
4.2.   Evaluation
First, we finetune a model pre-finetuned on three different multi-task sets on the BABE task and compare it against the single-task RoBERTa baseline. The multi-task sets are the following:
  • MTL:Random - Random Subset of 10 tasks
  • MTL:GradTS - Subset of 10 tasks selected by GradTS algorithm
  • MTL:All - Set of all tasks
We also evaluate the model pre-finetuned on the set of all tasks on MBIB. We follow the guidelines set by Wessel et al. (2023) for the evaluation. Given that MAGPIE’s pre-training data includes portions of the MBIB data, we ensure that the test set for each task in MBIB is not exposed to the model during its training or validation phases.
Figure 4: Final F1 score on a test set averaged over 30 random seeds. All three MTL approaches outperform baseline STL finetuning. Pre-finetuning on all LBM tasks results in significantly improved performance.
Multi-Task Learning Performance. Section 4.2 summarizes our performance results on the BABE dataset. We observe that all of our MTL pre-finetuning schemas lead to performance improvements. In particular, pre-finetuning on all tasks from LBM yields a SOTA performance on the BABE dataset, achieving an 84.1% F1 score and a relative improvement of 3.3% compared to the previous baseline by Spinde et al. (2021c). While both MTL baselines - Muppet Aghajanyan et al. (2021) and UnifiedM2 Lee et al. (2021b) outperform single-task baseline, they underperform all of our MTL models.
Task scaling. GradTS task selection outperforms random tasks on average performance, yet our experiment suggests task number scaling is more crucial. This is consistent with Muppet and ExT5 results Aghajanyan et al. (2021); Aribandi et al. (2021b), indicating MTL can compensate for scarce high-quality media bias datasets through general bias representation from other tasks. It also supports Kirstein et al. (2022)’s finding that sufficient related tasks can substitute the original task.
Model F1 Acc loss
Baseline (RoBERTa base) 80.83 (±0.69) 81.19 (±0.69) 43.6 (±3.54)
DA-RoBERTa 77.83 (±1.4) 78.56 (±1.3) 47.84 (±2.97)
MUPPET 80.56 (±1.3) 81.18 (±1.16) 44.19 (±4.65)
UnifiedM2 81.91 (±0.91) 82.41 (±0.88) 44.86 (±3.99)
MTL:Random 81.88 (±1.02) 82.28 (±0.97) 40.35 (±1.73)
MTL:GradTS 82.32 (±0.79) 82.64 (±0.8) 40.96 (±2.36)
MTL:All 84.1 (±1.33) 84.44 (±1.25) 39.46 (±2.41)
Table 2: Performance of two MTL baseline models (Muppet, UnifiedM2) two single-task baslines (RoBERTa and DA-RoBERTa) and our three MTL models, on fine-tuning on BABE dataset and evaluating on the held-out test set. The results are averaged over 30 random seeds.
On MBIB, our performance surpasses that of the five models evaluated by Wessel et al. (2023) in 5 out of 8 tasks. Additionally, it improves upon the current baseline with an average F1-Score of 0.7692 across all tasks. However, this improvement is only marginal. More importantly, MAGPIE outperforms the single-task RoBERTa baseline on MBIB, showing that an MTL approach is beneficial in media bias detection.
Figure 5: Evaluation F1 score during the final finetuning where MTL: All shows superior performance in training-step efficiency. The values are averaged over 30 random seeds. The bands mark the lowest and highest values.
Step efficiency. In addition to the performance improvements achieved, we also assess our model training efficiency. In Figure 5, we show the F1 score on the development set for the BABE task, averaged over all 30 runs. Our findings show that Multi-Task Learning only requires 1̃5% of the training steps used in single-task finetuning on BABE. This result demonstrates the high training-step efficiency of MAGPIE in media bias classification, making MTL implementations in the media bias domain more viable in the future.
4.3.   LBM taxonomy analysis
In Section 3.1, following Aribandi et al. (2021b), we introduce data task families. Aribandi et al. (2021b) uses task families for selection and knowledge transfer. To assess task families’ significance in LBM taxonomy, we train each pair of families together, investigating knowledge transfer.
To account for potential negative transfer within families, we first calculate the average transfer within each family and use it as a baseline for measuring transfer between families. We train tasks from the same family together and report the average change in task performance, as depicted in Figure 6. Negative knowledge transfer is prevalent across most of our task families. However, we observe two exceptions: the hate-speech and stance detection families, where multi-task training leads to an average improvement in performance.
Figure 6: Average performance change per task family. Stance detection and hate speech are the only families, on average, benefitting from Multi-Task Learning.
Next, we measure knowledge transfer between families by training each pair together. We report the average impact of each family on others and the average benefit each family gets through training with others, summarized in Table 3.
Our results show that, on average, the Emotionality and Sentiment analysis families provide positive transfer learning to other families. Conversely, we observe that the Fake News family benefits from knowledge transfer from every other family, with an average improvement of 1.7%. On the other hand, the Emotionality family is significantly impaired by negative transfer from other task families.
Figure 7: Average performance change when training tasks together. The change is measured with respect to transfer within each respective family (see results in section Figure 6.). The values in the horizontal axis represent the received transfer by the family on the y-axis. E.g., when training Emotionality and Subjective bias together, Emotionality gets worse by 11% whereas Subjectivity improves by 1.5%.
The full table of transfers can be found in Figure 7. Considering that only two families show positive transfer learning, with marginal effects of 0.11% and 0.34%, we conclude that the task families used in the construction of LBM are generally unsuitable for effectively utilizing knowledge transfer. We discuss this again in Section 5.
Task Family Transfer from Transfer to
media bias -2.07% -0.94%
subjectivity -1.26% 0.89%
hate speech -0.87% 0.17%
gender bias -1.01% -1.07%
sentiment analysis 0.11% 0.72%
fake news -0.13% 1.79%
group bias -1.04% 0.09%
emotionality 0.34% -6.56%
stance detection -0.79% -1.83%
Table 3: Evaluation of averaged transfer between task families.
4.4.   Resurrection and HSES evaluation
To evaluate the Resurrection and HSES methods in combination with other training strategies, we run a grid search on the following training strategies: HSES, Resurrection, Loss Scaling and Gradient Aggregation. We calculate the average evaluation loss for both Resurrection and HSES methods across 20 tasks randomly selected from the LBM collection. The boxplot in Figure 8 shows that both methods reduce the loss by 5% and decrease the variance across different training setups by 85%. However, we hypothesize that a random constellation of tasks [^9] Particularly variance in task sizes and quality. can have a non-trivial effect on the evaluation of our technique; thus, we opt for robust examination of the methods in future work.
(a) Resurrection
(b) HSES
Figure 8: An averaged evaluation loss of 20 tasks when trained with different training strategies. Both Resurrection and HSES approaches, compared to the vanilla setting, lead to significantly lower variance and overall lower loss.
5.   Conclusion
This paper contributes to media bias detection by the development of MAGPIE, our large-scale multi-task learning (MTL) approach that enhances traditional models with a multi-task structure.
Additionally, we present Large Bias Mixture (LBM), a compilation of 59 bias-related tasks. This broad collection serves as a resource for the pre-training of MAGPIE. To the best of our knowledge, it is the first available MTL resource tailored specifically towards media bias.
Our study investigates the dynamics of transfer learning across tasks and task families within media bias. Despite the occurrence of negative transfer among several tasks, scaling the pre-training setup to all collected tasks in Multi-Task Learning results in a 3.3% improvement over the previous state-of-the-art, making it the biggest advancement in neural media bias classification so far. Furthermore, we report that finetuning MAGPIE on the BABE dataset only requires 15% of steps compared to RoBERTa single-task approaches. These findings underscore the effectiveness and potency of Multi-Task Learning in highly specific classification domains such as media bias.
While results suggest benefits in scaling tasks, we see more promise in novel tasks rooted in media bias, suggesting deeper exploration over simply expanding the task spectrum. Understanding families and tasks in datasets necessitates systematic analysis of label definitions, rater agreement, and inter-relatedness of dataset creation strategies. As media bias is emerging globally, incorporating multilingual models is a natural extension.
Limitations
We acknowledge the necessity for a more comprehensive analysis of the performance of the HSES and Resurrection methods to ensure robust evaluation.
Given the significant computational resources required for a single multi-task training of all tasks, we allocated these resources towards robustly evaluating the model performance rather than conducting an in-depth analysis of the optimization techniques. Consequently, the presented methods may be less reliable and have limited applicability.
While there are various models available, previous research suggests that RoBERTa exhibits strong performance on media bias tasks Wessel et al. (2023); Spinde et al. (2021c). However, due to resource constraints, we were unable to explore models with different architectures and further refine our selection.
As the landscape of publicly available high-quality datasets for media bias is not as extensive as desired, we acknowledge our inability to capture all manifestations of media bias comprehensively. As mentioned in Section 5, a systematic and comprehensive analysis of the dataset landscape will be part of our future work.
Furthermore, although analyzing media bias on a sentence level enables a detailed examination of occurring biases, certain forms of bias, such as bias resulting from the omission of information, extend beyond the linguistic aspects of individual statements. Detecting such biases may require considering an external context.
We leave it to future work to investigate other levels of granularity, such as paragraph or article levels. In addition to these technical limitations, conducting a detailed analysis of agreement and label distribution in all utilized datasets will be necessary to make stronger claims about which datasets provide more reliable coverage of the overall concept of media bias. This is particularly important since media bias is a complex phenomenon that is not always easily identified during the annotation-gathering process.
Ethics Statement
Detecting (and as a result also highlighting) media bias instances can mitigate the negative effects of media bias, e.g., on collective decision-making Baumer et al. (2015). However, media bias strongly depends on the context and individuals involved, making it a sensitive issue. Some bias forms depend on factors other than the content, e.g., a different text perception due to a reader’s background. Therefore, a public classification of possible polarization and one-sidedness in the news must be performed transparently and reliably. The datasets used for training the classifier must be transparent and, ideally, constantly monitored by independent experts. To address possible implications of MAGPIE or related media bias applications cautiously, it is crucial to constantly control the classifications and datasets with recent findings on the perception of media bias, which is a main part of our ongoing and future work. To do so, we use resources such as, e.g., standardized questionnaires on media bias perception Spinde et al. (2021a).
When bias detection is balanced and transparent, it can positively affect collective opinion formation and decision-making. We see no immediate negative ethical or societal impacts of our work. However, in addition to system transparency, we want to highlight that we believe it to be required for future datasets to report in greater detail about a manipulation protection strategy when developing, training, and presenting any media bias classifier. To ensure the validity of media bias detection systems, it is essential to prevent participants, especially in public studies, from manipulating algorithms by, for example, flagging neutral content as biased. Therefore, annotations should be compared among multiple users to ensure a higher level of trustworthiness. Most of the datasets available report only limitedly about such strategies. In open or crowdsourcing scenarios, it is important to collect user characteristics and deliberately include specific content that aims to provide obvious answers but may be answered differently when users follow certain patterns. This approach helps in detecting and mitigating potential biases introduced by participants, thereby maintaining the integrity and reliability of the media bias detection process. To ensure and propose stronger standards in the future, we aim to analyze all LBM datasets with regard to potential inherent bias in future work.
Acknowledgments
We are grateful for the financial support of this project provided by the Hanns-Seidel Foundation, the DAAD (German Academic Exchange Service, program number 57515303 and 57515245), the Lower Saxony Ministry of Science and Culture, the VW Foundation and the AI Center of Czech Technical University. Furthermore, the authors would like to express their gratitude towards Jan Drchal for administrative support.
6.   Bibliographical References
[…]
Task Family Dataset # sentences Task Subjective bias SUBJ Pang and Lee (2004) 10.000 Binary Classification Wiki Neutrality Corpus Pryzant et al. (2020) 52.036 Token-level Classification NewsWCL50 Hamborg et al. (2019) 731 Regression CW_HARD Hube and Fetahu (2019) 6.843 Binary Classification News bias MultiDimNews Färber et al. (2020) 2.015 Multi-Label Classification BASIL Fan et al. (2019) 7.987 Multi-Class Classification Token-Level Classification Starbucks Lim et al. (2020) 866 Regression SemEval2023Task3 Piskorski et al. (2023) 5.219 Binary Classification BABE Spinde et al. (2021c) 3.672 Binary Classification Token-Level Classification Hate speech OffensiveLanguage Davidson et al. (2017) 23.198 Multi-Class Classification OnlineHarassmentDataset Golbeck et al. (2017) 19.613 Binary Classification WikiDetoxToxicity Wulczyn et al. (2017a) 138.827 Regression WikiDetoxAggression Wulczyn et al. (2017a) 101.159 Binary Classification Regression Jigsaw AI (2019) 101.060 Binary Classification MeTooMA Gautam et al. (2020) 7.388 Multi-Label Classification WikiMadlibs Dixon et al. (2018) 74.972 Binary Classification HateXplain Mathew et al. (2021) 18.962 Multi-Class Classification Multi-Label Classification Token-Level Classification HateSpeechTwitter Founta et al. (2018) 48.572 Multi-Class Classification Gender bias GAP Webster et al. (2018) 4.373 Multi-Class Classification RtGender Voigt et al. (2018a) 21.690 Binary Classification MDGender Dinan et al. (2020) 2.332 Multi-Class Classification TRAC2 Safi Samghabadi et al. (2020) 3.983 Binary Classification Funpedia Miller et al. (2017) 11.256 Multi-Class Classification WizardsOfWikipedia Dinan et al. (2019) 29.777 Multi-Class Classification Sentiment analysis SST2 Socher et al. (2013a) 9.436 Binary Classification IMDB Maas et al. (2011a) 13.139 Binary Classification MPQA Wilson (2008) 3.508 Binary Classification SemEval2014 Pontiki et al. (2014) 5.794 Token-Level Classification AmazonReviews Zhang et al. (2015b) 167.396 Binary Classification Fake news LIAR Wang (2017a) 12.742 Regression FakeNewsNet Shu et al. (2020a) 21.299 Binary Classification PHEME Kochkina et al. (2018) 5.022 Binary Classification Multi-Class Classification Emotionality GoodNewsEveryone Bostan et al. (2020) 4.428 Token-Level Classification Token-Level Classification BU-NEMO Reardon et al. (2022) 12.576 Multi-Class Classification EmotionTweets Krommyda et al. (2021b) 195.744 Multi-Class Classification DebateEffects Sridhar and Getoor (2019) 6.941 Regression Group bias CrowSPairs Nangia et al. (2020) 3.009 Binary Classification Multi-Class Classification Token-Level Classification StereoSet Nadeem et al. (2021b) 4.170 Binary Classification Multi-Class Classification StereotypeDataset Pujari et al. (2022) 2.208 Binary Classification Multi-Label Classification RedditBias Barikeri et al. (2021b) 10.395 Binary Classification Multi-Class Classification Token-Level Classification Stance detection SemEval2023Task4 Mirzakhmedova et al. (2023) 5.219 Binary Classification VaccineLies Weinzierl and Harabagiu (2022) 4.497 Multi-Class Classification SemEval2016Task6 Mohammad et al. (2016) 4.849 Multi-Class Classification WTWT Conforti et al. (2020) 24.681 Multi-Class Classification MultiTargetStance Sobhani et al. (2017) 4.430 Multi-Class Classification GWSD Luo et al. (2020) 2.010 Multi-Class Classification $\sum$ 1.210.084
Table 4: References and description to all 59 Tasks (46 datasets) in LBM collection.